# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix C $-\ell$ Converse Analysis for Bernoulli Designs
Therefore, for a specific $k+1\leq j^{\prime}\leq n$, conditioning on $(\mathbf{X}_{s},\mathbf{Y})$, by the randomness of the $j^{\prime}$-th column of $\mathbf{X}$ (i.e., the placement of the $j^{\prime}$-th item), we have
$G_{j,j^{\prime},1}\sim\text{Bin}\Big{(}\rho ne^{-\nu}(1+o(1)),\frac{\nu}{k}\Big{)},$ (62)
$G_{j,j^{\prime},2}\sim\text{Bin}\Big{(}(1-\rho)ne^{-\nu}(1+o(1)),\frac{\nu}{k}\Big{)}.$ (63)
@@
In order to establish a lower bound on the probability of (37), we introduce a parameter $d$ (to be chosen later) satisfying
$d>\max\Big{\{}\frac{C(1-2\zeta)}{\rho},0\Big{\}}$ (64)
and consider the event $G_{j,j^{\prime},1}=\frac{d\rho\nu e^{-\nu}n}{k}$ (assumed to be positive integer). We then use (25) from Lemma 1 to obtain
$\mathbb{P}\left(G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)$ (65)
$\geq\mathbb{P}\left(G_{j,j^{\prime},1}=\frac{d\rho\nu e^{-\nu}n}{k}\right)\mathbb{P}\left(G_{j,j^{\prime},2}<\frac{n\nu e^{-\nu}}{k}\big{(}d\rho-(1-2\zeta)C\big{)}\right)$ (66)
$\geq\Theta\Big{(}\frac{1}{\sqrt{n/k}}\Big{)}\exp\left(-\rho ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{d(1+o(1))\nu}{k}\big{\|}\frac{\nu}{k}\Big{)}\right)$ (67)
@@
×$\quad\times\Theta\Big{(}\frac{1}{\sqrt{n/k}}\Big{)}\exp\left(-(1-\rho)ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{(d\rho-(1-2\zeta)C)(1+o(1))}{1-\rho}\frac{\nu}{k}\big{\|}\frac{\nu}{k}\Big{)}\right)$ (68)
$=\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{[}g(C,\zeta,d,\rho)+o(1)\big{]}\right),$ (69)
@@
where in (69) we define $g(C,\zeta,d,\rho)$ that depends on $(C,\zeta)$ only through $A:=C(1-2\zeta)$ as follows:
$g(C,\zeta,d,\rho)=\rho d\log d+\big{(}\rho d-A\big{)}\log\Big{(}\frac{\rho d-A}{1-\rho}\Big{)}+1-2\rho d+A,$ (70)
and (69) itself follows from $D(\frac{a\nu}{k}\|\frac{\nu}{k})=\frac{\nu}{k}(a\log a-a+1+o(1))$ and writing the pre-factors $\Theta((\frac{n}{k})^{-1/2})$ as $\exp(\frac{o(1)n}{k})$. Differentiating $g(C,\zeta,d,\rho)$ with respect to $d$, we obtain
$\frac{\partial g}{\partial d}=\rho\log\Big{(}\frac{d(\rho d-A)}{1-\rho}\Big{)},$ (71)
@@
which we observe is zero if and only if $\rho d^{2}-Ad-(1-\rho)=0$. Thus, setting $\frac{\partial g}{\partial d}=0$, we readily obtain that
$\min_{d>\max\{0,\frac{A}{\rho}\}}\leavevmode\nobreak\ g(C,\zeta,d,\rho)=g(C,\zeta,d^{*},\rho):=f_{2}(C,\zeta,\rho),$ (72)
where $\text{where }d^{*}=\frac{A+\sqrt{A^{2}+4\rho(1-\rho)}}{2\rho}.$ (73)
@@
Therefore, we let $d=d^{*}+o(1)$ (here the $o(1)$ term only serves to ensure that $\frac{d\rho\nu e^{-\nu}n}{k}$ is a positive integer) to get from (69) the strongest bound:
$\mathbb{P}\left(G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)\geq\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\right).$ (74)
@@
Recall that to ensure the failure of MLE, with a chosen $j\in\mathcal{K}_{C,\zeta}$ at hand, the condition (37) should hold for some $k+1\leq j^{\prime}\leq p$. The placements of the different non-defective items are independent, and by implicitly conditioning on $(\mathbf{X}_{s},\mathbf{Y})$ we have
@@
$\mathbb{P}\left(\exists\leavevmode\nobreak\ k+1\leq j\leq p,\text{ s.t. }G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)$ (75)
@@
$\quad=1-\mathbb{P}\left(\forall\leavevmode\nobreak\ k+1\leq j\leq p,\leavevmode\nobreak\ G_{j,j^{\prime},1}-G_{j,j^{\prime},2}\leq(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)$ (76)
$\quad=1-\left[\mathbb{P}\left(G_{j,j^{\prime},1}-G_{j,j^{\prime},2}\leq(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)\right]^{p-k}$ (77)
$\quad\geq 1-\left[1-\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\right)\right]^{p-k}.$ (78)
Moreover, using $1-a\leq e^{-a}$ we have
$\left[1-\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\right)\right]^{p-k}\leq\exp\left(-(p-k)\exp\Big{(}-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}\right).$ (79)
@@
Thus, to ensure that (78) behaves as $1-o(1)$, it suffices to have $\log(p-k)-\frac{n\nu e^{-\nu}}{k}(f_{2}(C,\zeta,\rho)+o(1))\to\infty$. Substituting $\log(p-k)\sim\log p\sim\frac{1}{1-\theta}\log\frac{p}{k}$, we obtain that if
$\frac{1}{1-\theta}\log\frac{p}{k}-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\geq\eta_{2}\log\frac{p}{k}$ (80)
holds for some $\eta_{2}>0$, then (C2) holds with $1-o(1)$ probability. Note that the above condition can also be written as
$n\leq\frac{(1-\eta_{2})\frac{1}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}}$ (81)
for some $\eta_{2}>0$ possibly different from the one in (80).
@@
By merging $\eta_{1},\leavevmode\nobreak\ \eta_{2}$ and $o(1)$ into a factor of $1-\eta$, the statement “(58) and (81) simultaneously hold” can be written as
$n\leq(1-\eta)\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\}}$ (82)
for some $\eta>0$. Under this condition, the optimal decoder MLE fails with $1-o(1)$ probability. Further optimizing over $(C,\zeta)$ to render the tightest converse bound, we arrive at the threshold
$\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C>0,\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\}},$ (83)
where $f_{1}(C,\zeta,\rho)$ is given in (56), and $f_{2}(C,\zeta,\rho)$ is defined in (72) and (70). This establishes the second term in (4).
@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
Recall that in this part of the analysis, we restrict our attention to $\ell\in[1,\frac{k}{\log k}]$. Our strategy is to show that there is $o(1)$ probability of the restricted MLE decoder (30) failing. This is significantly more involved than the converse analysis because we need to consider the failure events associated with all $\ell\in[1,\frac{k}{\log k}]$ and the corresponding feasible $(C,\zeta)$ pairs, and finally bound the overall failure probability. Due to the symmetry of the design with respect to re-ordering items, it suffices to consider the fixed defective set $s=[k]$, meaning that items $1,2,\dotsc,k$ are defective and items $k+1,k+2,\dotsc,p$ are non-defective.
@@
For $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$ we will use the notations $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and their corresponding cardinalities (e.g., $M_{\mathcal{J}}=|\mathcal{M}_{\mathcal{J}}|$, $N_{0}=|\mathcal{N}_{0}|$) from Lemma 3. Conditioned on $\mathbf{X}_{s}$ and $\mathbf{Y}$, these index sets and their cardinalities are deterministic. Given $\ell\in[1,\frac{k}{\log k}]$, we say $(C,\zeta)\in[0,\infty)\times[0,1]$ is feasible if $\frac{Cn\nu e^{-\nu}\ell}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}$ are integers, subject to the restriction that $(C,\zeta)=(0,0)$ is the only feasible pair with $C=0$. Given feasible $(C,\zeta)$, we define
@@
as in (32). For given $(\mathcal{J},\mathcal{J}^{\prime})\in\mathcal{K}_{\ell,C,\zeta}\times([p]\setminus s)$ with $|\mathcal{J}^{\prime}|=\ell$, recall that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ that contain some item from $\mathcal{J}^{\prime}$, and ${G}_{\mathcal{J},\mathcal{J}^{\prime},2}$ denotes the number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$ that contain some item from $\mathcal{J}^{\prime}$. All of these definitions are consistent with Lemma 3.
To bound the probability that restricted MLE fails, we first invoke Lemma 3(b) to obtain a necessary condition for the failure of restricted MLE. In particular, Lemma 3(b) implies the following: If restricted MLE fails and returns some $s^{\prime}$ satisfying $|s\setminus s^{\prime}|\in[1,\frac{k}{\log k}]$, then there exist some feasible $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to the specific $\ell:=|s\setminus s^{\prime}|$, such that (C1) and (C2) below simultaneously hold:
@@
    $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}.$ (84)
@@
In general, we can have integer-valued $\frac{Cn\nu e^{-\nu}\ell}{k}$ with arbitrarily large $C$. In this part, we first address the cases of $C>C_{0}$ for some sufficiently large absolute constant $C_{0}$, so that we can simply concentrate on $C=O(1)$ later. For specific $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$, we let the probability of a test belonging to $\mathcal{M}_{\mathcal{J}}$ be
$P_{1}:=\Big{(}1-\big{(}1-\frac{\nu}{k}\big{)}^{\ell}\Big{)}\Big{(}1-\frac{\nu}{k}\Big{)}^{k-\ell}=\frac{\ell\nu e^{-\nu}(1+o(1))}{k},$ (85)
@@
((C1) and (C2) hold for some feasible $\mathbb{P}\Big{(}\textbf{(C1)}\text{ and }\textbf{(C2)}\text{ hold for some feasible }(C,\zeta)\in(C_{0},\infty)\times[0,1]\Big{)}$ (86)
$\leq\mathbb{P}\Big{(}\mathcal{K}_{\ell,C,\zeta}\neq\varnothing\text{ holds for some feasible }(C,\zeta)\in(C_{0},\infty)\times[0,1]\Big{)}$ (87)
$\leq\mathbb{P}\Big{(}\exists\leavevmode\nobreak\ \mathcal{J}\subset s\text{ with }|\mathcal{J}|=\ell,\leavevmode\nobreak\ \text{s.t. }M_{\mathcal{J}}>\frac{C_{0}n\nu e^{-\nu}\ell}{k}\Big{)}$ (88)
$\leq\binom{k}{\ell}\mathbb{P}\left(\text{Bin}\Big{(}n,\frac{\ell\nu e^{-\nu}(1+o(1))}{k}\Big{)}>\frac{C_{0}n\nu e^{-\nu}\ell}{k}\right)$ (89)
$\leq\exp\left((1+o(1))\ell\log k-\frac{\ell n}{k}\nu e^{-\nu}\big{(}C_{0}\log C_{0}-C_{0}+1+o(1)\big{)}\right)$ (90)
$\leq k^{-5},$ (91)
where in (87) we get an upper bound by only considering (C1) and not (C2), in (88) we further relax $\mathcal{K}_{\ell,C,\zeta}\neq\varnothing$ to the first condition of $\mathcal{K}_{\ell,C,\zeta}$ (namely $M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}$; see (32)), in (89) we apply a union bound over all $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$, in (90) we use the Chernoff bound (see (24) in Lemma 1), and (91) holds by letting $C_{0}=\Theta(1)$ be sufficiently large (recall that we have $\frac{n}{k}=\Theta(\log k)$). Therefore, for a given $\ell\in[1,\frac{k}{\log k}]$, (C1) and (C2) simultaneously hold for some feasible $(C,\zeta)\in(C_{0},\infty)\times[0,1]$ with probability no more than $k^{-5}$.

### 23 Oct 2024 v2

@@ Appendix C $-\ell$ Converse Analysis for Bernoulli Designs
For a specific non-defective $j^{\prime}\in\{k+1,\dotsc,n\}$, conditioning on $(\mathbf{X}_{s},\mathbf{Y})$ satisfying $\mathscr{A}_{1}$, the randomness of the $j^{\prime}$-th column of $\mathbf{X}$ (i.e., the placements of the $j^{\prime}$-th item) gives
$G_{j,j^{\prime},1}\sim\text{Bin}\Big{(}\rho ne^{-\nu}(1+o(1)),\frac{\nu}{k}\Big{)},$ (C.26)
$G_{j,j^{\prime},2}\sim\text{Bin}\Big{(}(1-\rho)ne^{-\nu}(1+o(1)),\frac{\nu}{k}\Big{)}.$ (C.27)
@@
In order to establish a lower bound on the probability of the “bad event” (C.1), we introduce a parameter $d$ (to be chosen later) satisfying
$d>\max\Big{\{}\frac{C(1-2\zeta)}{\rho},0\Big{\}}$ (C.28)
and consider the event $G_{j,j^{\prime},1}=\frac{d\rho\nu e^{-\nu}n}{k}$ (assumed to be positive integer). Thus, $d$ parametrizes the number of “relevant” positive tests that $j^{\prime}$ is placed in. The idea is that while the condition (C.1) could hold for a wide range of $d$ values, it suffices for our purposes to identify a single choice that dominates the error probability.
Accordingly, we use the anti-concentration bound (A.3) from Lemma A.1 to obtain the following:
$\mathbb{P}\left(G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)$ (C.29)
$\geq\mathbb{P}\left(G_{j,j^{\prime},1}=\frac{d\rho\nu e^{-\nu}n}{k}\right)\mathbb{P}\left(G_{j,j^{\prime},2}<\frac{n\nu e^{-\nu}}{k}\big{(}d\rho-(1-2\zeta)C\big{)}\right)$ (C.30)
$\geq\Theta\Big{(}\frac{1}{\sqrt{n/k}}\Big{)}\exp\left(-\rho ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{d(1+o(1))\nu}{k}\big{\|}\frac{\nu}{k}\Big{)}\right)$ (C.31)
@@
×$\quad\times\Theta\Big{(}\frac{1}{\sqrt{n/k}}\Big{)}\exp\left(-(1-\rho)ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{(d\rho-(1-2\zeta)C)(1+o(1))}{1-\rho}\frac{\nu}{k}\big{\|}\frac{\nu}{k}\Big{)}\right)$ (C.32)
$=\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{[}g(C,\zeta,d,\rho)+o(1)\big{]}\right),$ (C.33)
@@
where in (C.33) we define $g(C,\zeta,d,\rho)$ that depends on $(C,\zeta)$ only through $A:=C(1-2\zeta)$ as follows:
$g(C,\zeta,d,\rho)=\rho d\log d+\big{(}\rho d-A\big{)}\log\Big{(}\frac{\rho d-A}{1-\rho}\Big{)}+1-2\rho d+A,$ (C.34)
and (C.33) itself follows from $D(\frac{a\nu}{k}\|\frac{\nu}{k})=\frac{\nu}{k}(a\log a-a+1+o(1))$ and writing the pre-factors $\Theta((\frac{n}{k})^{-1/2})$ as $\exp(\frac{o(1)n}{k})$.
Next, we seek to find the choice of $d$ that gives the strongest bound. Differentiating $g(C,\zeta,d,\rho)$ with respect to $d$, we obtain
$\frac{\partial g}{\partial d}=\rho\log\Big{(}\frac{d(\rho d-A)}{1-\rho}\Big{)},$ (C.35)
@@
which is zero if and only if $\rho d^{2}-Ad-(1-\rho)=0$. Thus, setting $\frac{\partial g}{\partial d}=0$, we readily obtain that
$\min_{d>\max\{0,\frac{A}{\rho}\}}\leavevmode\nobreak\ g(C,\zeta,d,\rho)=g(C,\zeta,d^{*},\rho):=f_{2}(C,\zeta,\rho),$ (C.36)
where $\text{where }d^{*}=\frac{A+\sqrt{A^{2}+4\rho(1-\rho)}}{2\rho}.$ (C.37)
@@
Therefore, we let $d=d^{*}+o(1)$ (here the $o(1)$ term only serves to ensure that $\frac{d\rho\nu e^{-\nu}n}{k}$ is a positive integer) to get from (C.33) the strongest bound:
$\mathbb{P}\left(G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)\geq\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\right).$ (C.38)
@@
Recall that to ensure the failure of MLE, with a chosen $j\in\mathcal{K}_{C,\zeta}$ at hand, the condition (C.1) should hold for some $k+1\leq j^{\prime}\leq p$. The placements of the different non-defective items are independent, and by implicitly conditioning on $(\mathbf{X}_{s},\mathbf{Y})$ we have
@@
$\mathbb{P}\left(\exists j\in\{k+1,\dotsc,p\}\text{ s.t. }G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)$ (C.39)
@@
$\quad=1-\mathbb{P}\left(\forall j\in\{k+1,\dotsc,p\},\leavevmode\nobreak\ G_{j,j^{\prime},1}-G_{j,j^{\prime},2}\leq(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)$ (C.40)
$\quad=1-\left[\mathbb{P}\left(G_{j,j^{\prime},1}-G_{j,j^{\prime},2}\leq(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}\right)\right]^{p-k}$ (C.41)
$\quad\geq 1-\left[1-\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\right)\right]^{p-k}.$ (C.42)
Moreover, using $1-a\leq e^{-a}$, we have
$\left[1-\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\right)\right]^{p-k}\leq\exp\left(-(p-k)\exp\Big{(}-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}\right).$ (C.43)
@@
Thus, to ensure that (C.42) behaves as $1-o(1)$, it suffices to have $\log(p-k)-\frac{n\nu e^{-\nu}}{k}(f_{2}(C,\zeta,\rho)+o(1))\to\infty$. Substituting $\log(p-k)\sim\log p\sim\frac{1}{1-\theta}\log\frac{p}{k}$, we obtain that if
$\frac{1}{1-\theta}\log\frac{p}{k}-\frac{n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\geq\eta^{\prime}_{2}\log\frac{p}{k}$ (C.44)
holds for some $\eta^{\prime}_{2}>0$, then (C2) holds with $1-o(1)$ probability. Re-arranging, we obtain the following condition on $n$ for arbitrarily small $\eta_{2}>0$:
$n\leq\frac{(1-\eta_{2})\frac{1}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}}.$ (C.45)
@@
By merging $\eta_{1},\leavevmode\nobreak\ \eta_{2}$ and $o(1)$ into a single parameter $\eta$, the statement “(C.22) and (C.45) simultaneously hold” can be written as
$n\leq(1-\eta)\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\}},$ (C.46)
where $\eta>0$ is arbitrarily small. Under this condition, the optimal decoder MLE fails with $1-o(1)$ probability. Further optimizing over $(C,\zeta)$ to render the tightest converse bound, we arrive at the threshold
$\frac{k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C>0,\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\}},$ (C.47)
where $f_{1}(C,\zeta,\rho)$ is given in (C.20), and $f_{2}(C,\zeta,\rho)$ is defined in (C.36) and (C.34). This establishes the second term in (1.4).
@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
Recall that in this part of the analysis, we restrict our attention to $\ell\in[1,\frac{k}{\log k}]$. Our strategy is to show that there is $o(1)$ probability of the restricted MLE decoder (B.2) failing. This is more involved than the converse analysis because we need to consider the failure events associated with all $\ell\in[1,\frac{k}{\log k}]$ and the corresponding feasible $(C,\zeta)$ pairs, and finally bound the overall failure probability. Due to the symmetry of the design with respect to re-ordering items, it suffices to again consider the fixed defective set $s=[k]$, meaning that items $1,2,\dotsc,k$ are defective and items $k+1,k+2,\dotsc,p$ are non-defective.
@@
For $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$, we will use the notations $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and their corresponding cardinalities (e.g., $M_{\mathcal{J}}=|\mathcal{M}_{\mathcal{J}}|$, $N_{0}=|\mathcal{N}_{0}|$) from Lemma B.1. Conditioned on $\mathbf{X}_{s}$ and $\mathbf{Y}$, these sets and their cardinalities are deterministic. Given $\ell\in[1,\frac{k}{\log k}]$, we say $(C,\zeta)\in[0,\infty)\times[0,1]$ is feasible if $\frac{Cn\nu e^{-\nu}\ell}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}$ are integers, subject to the restriction that $(C,\zeta)=(0,0)$ is the only feasible pair with $C=0$. Given a feasible pair $(C,\zeta)$, we define
@@
as in (B.4). For given $(\mathcal{J},\mathcal{J}^{\prime})\in\mathcal{K}_{\ell,C,\zeta}\times([p]\setminus s)$ with $|\mathcal{J}^{\prime}|=\ell$, recall that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ that contain some item from $\mathcal{J}^{\prime}$, and ${G}_{\mathcal{J},\mathcal{J}^{\prime},2}$ denotes the number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$ that contain some item from $\mathcal{J}^{\prime}$. All of these definitions are consistent with Lemma B.1.
To bound the probability that restricted MLE fails, we first invoke Lemma B.1(b) to obtain a necessary condition for failure. In particular, Lemma B.1(b) implies the following: If restricted MLE fails and returns some $s^{\prime}$ satisfying $|s\setminus s^{\prime}|\in[1,\frac{k}{\log k}]$, then there exists some feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to the specific $\ell:=|s\setminus s^{\prime}|$, such that (C1) and (C2) below simultaneously hold:
@@
  • (C2) There exist some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$ such that
    $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}.$ (D.1)
@@
In general, we can have integer-valued $\frac{Cn\nu e^{-\nu}\ell}{k}$ with arbitrarily large $C$. In this part, we first address the cases of $C>C_{0}$ for some sufficiently large absolute constant $C_{0}$, so that we can restrict to $C=O(1)$ later. For specific $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$, we let the probability of a test belonging to $\mathcal{M}_{\mathcal{J}}$ be
$P_{1}:=\Big{(}1-\big{(}1-\frac{\nu}{k}\big{)}^{\ell}\Big{)}\Big{(}1-\frac{\nu}{k}\Big{)}^{k-\ell}=\frac{\ell\nu e^{-\nu}(1+o(1))}{k},$ (D.2)
@@
((C1) and (C2) hold for some feasible $\mathbb{P}\Big{(}\textbf{(C1)}\text{ and }\textbf{(C2)}\text{ hold for some feasible }(C,\zeta)\in(C_{0},\infty)\times[0,1]\Big{)}$ (D.3)
$\leq\mathbb{P}\Big{(}\mathcal{K}_{\ell,C,\zeta}\neq\varnothing\text{ holds for some feasible }(C,\zeta)\in(C_{0},\infty)\times[0,1]\Big{)}$ (D.4)
$\leq\mathbb{P}\Big{(}\exists\leavevmode\nobreak\ \mathcal{J}\subset s\text{ with }|\mathcal{J}|=\ell,\leavevmode\nobreak\ \text{s.t. }M_{\mathcal{J}}>\frac{C_{0}n\nu e^{-\nu}\ell}{k}\Big{)}$ (D.5)
$\leq\binom{k}{\ell}\mathbb{P}\left(\text{Bin}\Big{(}n,\frac{\ell\nu e^{-\nu}(1+o(1))}{k}\Big{)}>\frac{C_{0}n\nu e^{-\nu}\ell}{k}\right)$ (D.6)
$\leq\exp\left((1+o(1))\ell\log k-\frac{\ell n}{k}\nu e^{-\nu}\big{(}C_{0}\log C_{0}-C_{0}+1+o(1)\big{)}\right)$ (D.7)
$\leq k^{-5},$ (D.8)
where in (D.4) we get an upper bound by only considering (C1) and not (C2), in (D.5) we further relax $\mathcal{K}_{\ell,C,\zeta}\neq\varnothing$ to the first condition of $\mathcal{K}_{\ell,C,\zeta}$ (namely $M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}$; see (B.4)), in (D.6) we apply a union bound over all $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$, in (D.7) we use the Chernoff bound (see (A.2) in Lemma A.1), and (D.8) holds by letting $C_{0}=\Theta(1)$ be sufficiently large (recall that we have $\frac{n}{k}=\Theta(\log k)$). Therefore, for a given $\ell\in[1,\frac{k}{\log k}]$, (C1) and (C2) simultaneously hold for some feasible $(C,\zeta)\in(C_{0},\infty)\times[0,1]$ with probability no higher than $k^{-5}$.